### v1

# SAFR: Neuron Redistribution for Interpretability

Abstract
Superposition refers to encoding representations of multiple features within a single neuron, which is common in transformers. This property allows neurons to combine and represent multiple features, enabling the model to capture intricate information and handle complex tasks. Despite promising performance, the model’s interpretability has been diminished. This paper presents a novel approach to enhance transformer interpretability by regularizing feature superposition. We introduce SAFR  [^1] SAFR: Superposition-Aware Feature Regularization, which simply applies regularizations to the loss function to promote monosemantic representations for important tokens while encouraging polysemanticity for correlated token pairs, where important tokens and correlated token pairs are identified via VMASK Chen and Ji (2020) and attention weights. With a transformer model on two classification tasks, SAFR improves interpretability without compromising prediction performance. Given an input to the model, SAFR provides an explanation by visualizing the neuron allocation and interaction within the MLP layers.
SAFR: Neuron Redistribution for Interpretability
Ruidi Chang  Chunyuan Deng  Hanjie Chen Department of Computer Science Rice University {ruidi.chang, hanjie}@rice.edu
1 Introduction
Individual neurons in neural networks can represent multiple features from the input. This phenomenon, known as superposition, improves the model’s ability to capture intricate relationships between features Olah et al. (2020) and complicates the understanding of the underlying processes behind the model’s decision-making Elhage et al. (2022). Facing these challenges, recent research like sparse autoencoders (SAEs) Huben et al. (2024) artificially decomposes the representation space into a sparse vector space through auxiliary networks. However, while SAEs provide a method to interpret features through combinations of sparse activations, there is still a lack of sufficient research on controlling neuron distribution.
In this paper, we ask the question: Can we enhance model interpretability by explicitly controlling the distribution of features across neurons? An intuitive approach is to encourage the development of monosemantic neurons Elhage et al. (2022); Bricken et al. (2023); Wang et al. (2024) by regulating activations. However, focusing solely on monosemanticity may limit the model’s ability to capture feature interactions, potentially hindering overall performance.
To address this challenge, we propose a novel method called Superposition-Aware Feature Regularization (SAFR) to enhance model interpretability by strategically reallocating neurons through a modified loss function, approached from two perspectives. As illustrated in Figure 1, our framework incorporates regularization techniques aimed at promoting monosemantic representations for important tokens, while simultaneously fostering polysemanticity among correlated token pairs. The identification of important tokens is achieved via a variational inference network, adapted from VMASK Chen and Ji (2020). Additionally, correlated token pairs are identified based on attention weights to foster polysemantic representation.
We evaluate SAFR on the SST-2 Socher et al. (2013) and IMDB Maas et al. (2011) datasets to assess its effectiveness. When the top 30% of words identified by SAFR are removed, we observe a significant drop in test accuracy—18.24% on SST-2 and 27.04% on IMDB. As we incrementally remove additional tokens, the accuracy continues to decline in a consistent manner. To further substantiate these findings, we visualize the neuron allocation within the FFN layers of a transformer block. Our experimental results confirm that SAFR can efficiently reallocate feature distributions across neurons while preserving a reasonable degree of polysemanticity, significantly improving interpretability in transformers while maintaining performance at comparable levels.
Figure 1: Basic structure of SAFR. i) Promote monosemanticity for important tokens after the embedding layers. ii) Leverage the attention mechanism to enhance polysemanticity among correlated token pairs.
2 Preliminaries
Given an input sequence of $n$ tokens $S=\langle s_{1},\ldots,s_{n}\rangle$ and a transformer $f(\cdot)$ with $L$ layers, let $h^{\ell}_{i}$ denote the hidden representation obtained at layer $\ell\in[1,\ldots,L]$ and position $i\in[1,\ldots,n]$. Following previous work  Scherlis et al. (2022); Elhage et al. (2022), we define interference, polysemanticity, and capacity as follows.
Interference
The interference ($I$) measures the overlap or similarity between the two representations at the same layer:
$I_{i,j}^{\ell}=\left(h^{\ell}_{i}\cdot h^{\ell}_{j}\right)^{2}$ (1)
Interference quantifies how much the hidden representations interfere with or influence each other. Higher interference means more overlap, suggesting the tokens share representational dimensions.
Polysemanticity
The polysemanticity ($P$) describes the extent to which a single hidden representation captures information from multiple tokens:
$P_{i}^{\ell}=\sum_{j\neq i}\left(\hat{h}^{\ell}_{i}\cdot h^{\ell}_{j}\right)^{2}$ (2)
Here, $\hat{h}^{\ell}_{i}$ denotes the normalized representation of $h^{\ell}_{i}$ (excluding its own magnitude). A high polysemanticity value indicates that a single dimension is “polysemantic”, meaning it represents information from multiple tokens simultaneously.
Capacity
The capacity ($C$) quantifies how much of a hidden representation is dedicated to representing token $i$:
$C_{i}^{\ell}=\frac{(h^{\ell}_{i}\cdot h^{\ell}_{i})^{2}}{\sum_{j}(h^{\ell}_{i}\cdot h^{\ell}_{j})^{2}}$ (3)
where $0\leq C_{i}\leq 1$ and $1\leq C\leq D$, $D$ denotes the hidden dimension size. A higher capacity value indicates that the representation at position $i$ is more focused on its own token.
3 Methodology
SAFR aims to enhance the interpretability of models by strategically reallocating neurons using superposition regularization strategy.
SST-2 IMDB
Model $\text{Acc}_{S}$ $\text{Acc}_{\tilde{S}^{(r)}}$ $\text{Acc}_{\tilde{S}^{(k)}}$ SRS $\text{Acc}_{S}$ $\text{Acc}_{\tilde{S}^{(r)}}$ $\text{Acc}_{\tilde{S}^{(k)}}$ SRS
Baseline 71.81 75.81 74.96 -3.15 80.88 74.31 76.92 3.96
$\lambda_{Imp}=0,\lambda_{Inter}=0$ 71.24 71.98 61.98 9.26 79.07 73.30 64.70 14.37
SAFR 72.96 75.3 54.72 18.24 78.23 74.52 51.19 27.04
Table 1: Evaluation on SST-2 and IMDB datasets, k = 30. $\text{Acc}_{S}$ denotes prediction accuracy for original test dataset. $\text{Acc}_{\tilde{S}^{(r)}}$ denotes prediction accuracy for dataset randomly drop kwords. $\text{Acc}_{\tilde{S}^{(k)}}$ denotes prediction accuracy for deleting kwords based on capacity. SAFR achieves a SRS score of 18.24 for SST-2 and 27.04 for IMDB, outperforming the baseline model, indicating an improvement on model interpretability.
The baseline is defined via original model in cross-entropy loss $\mathcal{L}_{\text{CE}}=-\frac{1}{M}\sum_{m=1}^{M}\sum_{g=1}^{G}y_{m}^{g}\log(\hat{y}_{m}^{g})$, where $M$ is the number of samples in the dataset and $G$ is the number of classes in text classification. To improve model efficiency and interpretability in a constrained space, we propose a two-part regularization: one promotes monosemantic representations for important tokens, while the other encourages polysemanticity for correlated token pairs to reduce redundancy.
The proposed loss function integrates these regularization terms accordingly:
$\mathcal{L}$ $=\mathcal{L}_{\text{CE}}+\lambda_{Imp}\cdot\mathcal{L}_{\text{Importance}}+\lambda_{Inter}\cdot\mathcal{L}_{\text{Interaction}}$
where $\lambda_{Imp}$ controls the importance loss term and $\lambda_{Inter}$ controls the interaction loss term.
Importance-Based Regularization
We apply the VMASK Chen and Ji (2020) between the embedding layer and the positional encoder to identify important tokens. VMASK Chen and Ji (2020) is a variational word mask layer that is inserted into a neural text classifier and trained with the model. It learns to limit the flow of globally irrelevant or noisy word-level feature information to subsequent network layers, thus forcing the model to focus on the important features for prediction. To encourage monosemanticity for important tokens, we introduce a regularization term $\mathcal{L}_{\text{Importance}}=\sum_{i=1}^{T}\sqrt{P_{i}^{V}/{E}}$, where $T$ is the maximum token length, $E$ represents the embedding dimension and $P_{i}^{V}$ denotes the polysemanticity for the hidden representation of the $i$-th token in the VMASK layer. This regularization term penalize important tokens with high polysemanticity.
Interaction-Based Regularization
We leverage the attention mechanism and employe the attention weights to identify correlated tokens. The $\alpha$-th self-attention head is described as follows:
$A_{\alpha}=\text{softmax}\left(\frac{Q_{\alpha}K_{\alpha}^{T}}{\sqrt{d_{k}}}\right)$
where $Q_{\alpha}=\mathbf{X}W^{Q}_{\alpha}$ and $\quad K_{\alpha}=\mathbf{X}W^{K}_{\alpha}$, with $Q_{\alpha},K_{\alpha}\in\mathbb{R}^{n\times d_{k}}$. $\mathbf{X}=E_{\text{pos}}(S^{\prime})$ denotes the input matrix to the attention layer, where $E_{\text{pos}}(S^{\prime})$ is the positional encoding applied to the output $S^{\prime}$ from the vmask layer. The score $A_{\alpha(i,j)}$ indicates how much attention token $i$ places on token $j$.
To encourage pairs of highly correlated features to exhibit high polysemanticity, we introduce a loss term $\mathcal{L}_{\text{Interaction}}=\sum_{\alpha}\sum_{i,j}\frac{1}{T^{2}}A_{\alpha(i,j)}\cdot(1-I_{i,j}^{A_{\alpha}})$, where $I_{i,j}^{A_{\alpha}}$ is the $Interference$ of the attention weights metric of the $\alpha$-th attention head. This loss term penalizes highly correlated features that exhibit low interference values.
Proposed Loss Function
The loss term is now defined as:
$\mathcal{L}$ $=\mathcal{L}_{\text{CE}}+\lambda_{Imp}\cdot\mathcal{L}_{\text{Importance}}+\lambda_{Inter}\cdot\mathcal{L}_{\text{Interaction}}$
$\!=\!-\!\sum_{m=1}^{M}\!\sum_{g=1}^{G}y_{m}^{g}\!\log(\hat{y}_{m}^{g})\!+\!\lambda_{Imp}\!\cdot\!\sum_{m=1}^{M}\!\sum_{i=1}^{T}\!\sqrt{P_{i}^{V}\!/\!E}$
$\quad+\lambda_{Inter}\cdot\sum_{m=1}^{M}\sum_{\alpha}\sum_{i,j}\frac{1}{T^{2}}A_{\alpha(i,j)}\!(1-I_{i,j}^{A_{\alpha}})$
4 Experiment Setup
The proposed method is evaluated on two classification tasks using a standard transformer model.
Datasets
We adopt 2 benchmark datasets: Stanford Sentiment Treebank binary version SST-2 Socher et al. (2013) and movie reviews IMDB Maas et al. (2011). Table 4 in Appendix A shows the statistics of the datasets.
Model
We use a typical transformer architecture with a single layer, following the standard setup Vaswani (2017). This includes the complete transformer framework with its attention mechanism and positional encodings. In the multi-layer perceptron (MLP) section, we use two feed forward sub-layers: we first expand the dimensionality by a factor of four, then reduce it back by the same factor, aligning with the commonly used configuration in transformer models. The model is using random embedding to avoid the influence of embedding information. Table 5 in Appendix A shows the statistics of the model.
Baseline
Since our objective is to investigate how regularization can modify neuron resource allocation to enhance interpretability, we employ a standard Transformer model, without any modifications, as the baseline for comparison.
Evaluation
To evaluate our method, we define an evaluation metric called Superposition Regularization Score (SRS). By deleting the top $k\%$ of tokens based on capacity, SRS calculates the average change in the prediction accuracy over all test data as follows:
$SRS(k)\!=\!\frac{1}{M}\!(\sum_{S=1}^{M}\!1\!\cdot\!(\hat{y}_{S}=y)\!-\!\sum_{S=1}^{M}\!1\!\cdot\!(\hat{y}_{\tilde{S}^{(k)}}\!=\!y))$
where $\tilde{S}^{(k)}$ is constructed by dropping the $k\%$ top-scored words from $S$. The SRS metric serves to evaluate how effectively the model arranges neurons to represent features. The SRS metric quantifies the alignment between neuron allocations and the semantic significance of tokens. By assessing the structure of neuron allocation, SRS provides insights into the interpretability of the model’s internal encoding of information. Higher SRS values indicate that the deleted words are important, therefore indicates a stronger superposition regularization.
5 Results and Analysis
SAFR Improves Interpretability.
Results in Table 1 demonstrate that SAFR achieves superior SRS scores of 18.24 and 27.04 for SST-2 and IMDB datasets respectively, surpassing the baseline model. This provides evidence for SAFR’s efficacy in improving interpretability. Notably, even without regularization ($\lambda_{Imp}=0$ and $\lambda_{Inter}=0$), the interpretability layer before the transformer block yields improved SRS scores, suggesting its influence on neuronal allocation. An ablation study of $\lambda_{Imp}$ and $\lambda_{Inter}$ is detailed in Table 2. When calibrating these regularization terms, two critical trade-offs emerge: Excessive emphasis $\lambda_{Imp}$ may compromise the model’s ability to capture feature interactions. Conversely, focusing solely on $\lambda_{Inter}$ without balancing feature importance can reduce attention to critical words. Additional ablation studies can be found in Appendix D.
SST-2
Model $\text{Acc}_{S}$ $\text{Acc}_{\tilde{S}^{(r)}}$ $\text{Acc}_{\tilde{S}^{(k)}}$ SRS
Baseline 71.81 75.81 74.96 -3.15
$\lambda_{Imp}=0$, $\lambda_{Inter}=0$ 71.24 71.98 61.98 9.26
$\lambda_{Imp}=0$, $\lambda_{Inter}=1$ 71.98 69.07 61.86 10.12
$\lambda_{Imp}=1$, $\lambda_{Inter}=0$ 71.47 72.16 55.35 16.12
$\lambda_{Imp}=1$, $\lambda_{Inter}=0.1$ 72.96 75.30 54.72 18.24
$\lambda_{Imp}=100$, $\lambda_{Inter}=100$ 61.46 62.04 51.06 10.40
Table 2: Ablation study of $\lambda_{Imp}$ and $\lambda_{Inter}$. $\lambda$ choosing influence prediction accuracy and model interpretability.
(a) Capacity Barchart of FFN1.
(b) Visualization of FFN1.
(c) Interference Heatmap of FFN1.
Figure 2: (a) Important words exhibit higher capacity. (b) Circle size represents feature capacity, with larger circles indicating greater capacity. Red lines denote positive correlations, blue lines indicate negative correlations, and shorter lines indicate stronger correlations. (c) Important words demonstrate lower polysemanticity, while correlated words pair exhibit relatively higher interference.
Sensitivity to $k$ Selection.
Figure 3 illustrates the effect of token removal based on capacity. The ktop-scored words are directly removed from the original text to ensure that the evaluation reflects the model’s ability to perform with reduced input information. As tokens are incrementally removed, the accuracy declines consistently. The greater decline observed in our model compared to the baseline suggests that our model is more interpretable.
Figure 3: Sensitivity to $k$ Selection. As tokens are incrementally removed, the accuracy declines consistently.
Average Capacity Across the Vocabulary.
Table 3 presents an analysis of the average capacity per token status (important or not w.r.t VMASK) on the entire test dataset. Tokens are classified based on VMASK scores, with the top 30% identified as important and the rest 70% as less important. These results demonstrate the effectiveness of SAFR in prioritizing and allocating greater representational capacity to features that are important for tasks.
Average Capacity SST-2 IMDB
All Tokens 0.1925 0.1372
Important Tokens 0.4141 0.2912
Less Important Tokens 0.0975 0.0712
Table 3: Average capacity metric for SST-2 and IMDB datasets. The metric reveals that important tokens exhibit significantly higher capacity scores compared to the overall average capacity.
This strategic allocation enhances the model’s ability to provide clearer and more meaningful insights into its decision-making processes, reinforcing its value as a tool for model interpretability.
Neuron Allocation Across Layers.
We conducted an analysis of neuron allocation across layers, with the visualizations presented in Appendix C. Our observations reveal that neurons within the embedding layer exhibit a random distribution, failing to effectively capture their relative importance or inter-neuronal interactions. The VMASK layer begins to identify important tokens at a global level, yet it lacks the capability to analyze token-to-token interactions. While the attention layer demonstrates proficiency in capturing inter-token relationships, the interpretation of token importance remains a debatable topic. The second feedforward neural network layer compresses information into a lower-dimensional space and tend to allocate neuron capacity in a uniform manner. The expansion of the feedforward neural network, denotes FFN1 layer in our model, facilitates the network’s ability to learn more intricate patterns and relationships within the data Elhage et al. (2021). Our oberservation reveals that important tokens exhibit greater monosemanticity, and correlated token pairs demonstrate higher interference, illustrated in Figure 2(a) and 2(c).
Visualization of FFN1
Figure 2(b) illustrates the enhanced interpretability achieved after applying SAFR. In the visualization, the size of the circles represents the capacity of individual features, where larger circles indicate greater capacity. The red and blue lines depict interference between features, with red lines corresponding to positive correlations and blue lines indicating negative correlations. The length of these lines reflects the strength of the correlation. This representation provides a clear and intuitive understanding of both the distribution of neuron and the relationships among features, showcasing the effectiveness of SAFR in improving the interpretability of FFN1 layer. By visualizing these dynamics, the figure highlights key insights into feature importance and their interactions within the model.
6 Related Work
Superposition in neural networks has gained attention, with foundational work by Arora et al. (2018); Goh (2016). Olah et al. (2020) developed this idea into the “superposition hypothesis” and initiated studies on mechanistic interpretability concerning polysemantic neurons and circuits. Lecomte et al. (2024) showed that polysemanticity can emerge incidentally through regularization and neural noise.
Elhage et al. (2022) illustrated superposition in simplified networks. Concurrently, theoretical explorations Scherlis et al. (2022); Henighan et al. (2023); Hänni et al. (2024); Gurnee et al. (2023); Marshall and Kirchner (2024); Hänni et al. (2024); Katta (2024) and applications in language models Chen et al. (2023, 2024) have expanded the field.
Interpretability efforts include works like Dreyer et al. (2024); Black et al. (2022); Wang et al. (2023). Challenges in lifelong knowledge editing Hu et al. (2024) and identifying universal feature spaces across models Lan et al. (2024) mark promising directions for future research.
7 Conclusion
In this work, we introduced SAFR, a novel approach to enhance model interpretability by strategically regularizing feature superposition. Experiments on SST-2 and IMDB show that SAFR improves interpretability, as measured by our SRS metric, without compromising prediction performance.
Our method provides insights into the relationship between superposition and interpretability in transformers and offers a framework for visualizing neuron allocation across layers. This work contributes to mechanistic interpretability and suggests promising directions for extending the approach to larger models and broader applications.
8 Limitation
This study has several limitations. First, experiments were conducted using a single-layer transformer model; future work should examine the scalability of SAFR with more complex architectures. Second, while focused on classification task, the applicability of SAFR to other NLP tasks—such as natural language inference, question answering, and text generation—remains unexplored. Third, there is a need for more comprehensive and standardized evaluation metrics to assess SAFR effectively. Finally, SAFR does not fully elucidate the causal mechanisms behind the model’s decision-making process. Addressing these challenges offers valuable opportunities for future research.
9 Ethic Statements
Our research focuses on understanding and controlling the inner workings of transformer models, without collecting or using any human data; no personal or sensitive information is handled in this study. All datasets used in this work are public.
Appendix A Dataset Statistics
This section provides the statistical summary for the datsets and model.
Datasets #Train #Dev #Test
SST2 6244 825 1749
IMDB 20k 5k 25k
Table 4: Summary statistics for the datasets, where # counts the number of examples in the train/dev/test sets.
Layer Dimension
Embedding (Input Dimension, 256)
FFN1 (256, 1024)
FFN2 (1024, 256)
Table 5: Summary statistic for the model: Layer dimensions.
Appendix B Cross Entropy Loss
Given an input sequence of $n$ tokens $S=\langle s_{1},\ldots,s_{n}\rangle$ and a transformer $f(\cdot)$ with $L$ layers, let $h^{\ell}_{i}$ denote the hidden representation obtained at layer $\ell\in[1,\ldots,L]$ and position $i\in[1,\ldots,n]$. Following previous work Scherlis et al. (2022); Elhage et al. (2022), we define interference, polysemanticity, and capacity as follows. The baseline is defined via the original model using the cross-entropy loss:
$\mathcal{L}_{\text{CE}}=-\frac{1}{M}\sum_{m=1}^{M}\sum_{g=1}^{G}y_{m}^{g}\log(\hat{y}_{m}^{g})$
where $M$ is the number of samples in the dataset and $G$ is the number of classes in text classification.
Additionally, for the input sequence $S$, let:
  • $\hat{y}_{t}$ represent the predicted probability distribution over the vocabulary at position $t$,
  • $y_{t}$ represent the true one-hot encoded distribution for the token at position $t$,
  • $V$ denote the size of the vocabulary, and
  • $c_{t}$ be the index of the correct token at position $t$.
The cross-entropy loss at position $t$ is defined as:
$\mathcal{L}_{t}=-\sum_{i=1}^{V}y_{t,i}\log(\hat{y}_{t,i})$
Since $y_{t}$ is a one-hot encoded vector, this simplifies to:
$\mathcal{L}_{t}=-\log(\hat{y}_{t,c_{t}})$
The total cross-entropy loss for the sequence $S$ is the average over all tokens in the sequence:
$\mathcal{L}=\frac{1}{n}\sum_{t=1}^{n}\mathcal{L}_{t}=-\frac{1}{n}\sum_{t=1}^{n}\log(\hat{y}_{t,c_{t}})$
where $\hat{y}_{t,c_{t}}$ is the predicted probability of the correct token at position $t$.
Appendix C Neuron Allocation Accross Layers
This section presents the observations regarding neuron allocation across the various layers.
Figure 4: Cross Layers Output-1: Capacity. Capacity for FFN1 is interpretable.
Figure 5: Cross Layers Output-2: Capacity. Capacity for FFN1 is interpretable.
Figure 6: Cross Layers Output: Interference. Interference for FFN1 is interpretable.
SST-2 IMDB
Model $\text{Acc}_{S}$ $\text{Acc}_{\tilde{S}^{(r)}}$ $\text{Acc}_{\tilde{S}^{(k)}}$ SRS $\text{Acc}_{S}$ $\text{Acc}_{\tilde{S}^{(r)}}$ $\text{Acc}_{\tilde{S}^{(k)}}$ SRS
Baseline 71.81 75.81 74.96 -3.15 80.88 74.31 76.92 3.96
$\lambda_{Imp}=0$, $\lambda_{Inter}=0$ 71.24 71.98 61.98 9.26 79.07 73.30 64.70 14.37
$\lambda_{Imp}=0$, $\lambda_{Inter}=0.01$ 71.98 72.04 57.29 14.69 78.38 74.12 61.55 16.83
$\lambda_{Imp}=0$, $\lambda_{Inter}=1$ 71.98 69.07 61.86 10.12 78.05 74.05 55.96 22.09
$\lambda_{Imp}=0$, $\lambda_{Inter}=100$ 62.66 78.56 81.25 -18.59 77.19 72.60 64.46 12.73
$\lambda_{Imp}=0.01$, $\lambda_{Inter}=0$ 73.47 69.70 59.35 14.12 78.63 74.48 50.42 28.21
$\lambda_{Imp}=0.01$, $\lambda_{Inter}=0.01$ 74.10 73.13 62.78 11.32 78.23 74.52 51.19 27.04
$\lambda_{Imp}=0.1$, $\lambda_{Inter}=0.1$ 72.78 70.67 60.43 12.35 77.64 74.06 50.05 27.59
$\lambda_{Imp}=0.1$, $\lambda_{Inter}=1$ 73.87 74.27 60.38 13.49 76.88 73.38 50.02 26.86
$\lambda_{Imp}=1$, $\lambda_{Inter}=0$ 71.47 72.16 55.35 16.12 74.93 71.57 50.00 24.93
$\lambda_{Imp}=1$, $\lambda_{Inter}=0.1$ 72.96 75.30 54.72 18.24 73.64 70.72 50.00 23.64
$\lambda_{Imp}=1$, $\lambda_{Inter}=1$ 71.58 73.70 56.89 14.69 74.10 70.89 50.00 24.10
$\lambda_{Imp}=10$, $\lambda_{Inter}=10$ 70.84 73.24 57.00 13.84 60.73 58.73 50.00 10.73
$\lambda_{Imp}=100$, $\lambda_{Inter}=0$ 59.86 64.95 51.57 8.29 52.04 51.81 50.09 1.95
$\lambda_{Imp}=100$, $\lambda_{Inter}=100$ 61.46 62.04 51.06 10.40 51.80 51.72 50.00 1.80
Table 6: Ablation Study on SST-2 and IMDB datasets. $\lambda$ choosing influence prediction accuracy and model interpretability.
Appendix D Regularization Hyperparameter Tuning
This section presents the results of hyperparameter tuning for $\lambda_{Imp}$ and $\lambda_{Inter}$, as summarized in Table 6.